OpenStudy (anonymous):

x+y-5=0 find the slope-intercept form

OpenStudy (anonymous):

x+y-5=0 +5 +5 --------------- x+y=5 -x -x -------------- y=-x+5 P.S. wrong sesction

OpenStudy (kohai):

Try posting it into math